A student measured a meter stick to be 150 gm. The student then placed a knife edge 

on 30-cm mark of the stick. If the student placed a 500-gm weight on 5-cm mark and 

a 300-gm weight on somewhere on the meter stick, the meter stick then was 

balanced. Where did the student place the 300-gram weight?



1
Expert's answer
2021-05-04T11:29:45-0400

Distance of 500 gm weight from pivot point,

"x_1 = 30 - 5 = 25 cm"


Distance of center of mass point,

"x_2 = 50 - 30 = 20 cm"


suppose 300 gm weight is put at distance r from 30cm mark,

"(500 \times 25) - (150 \times 20) - (300 \times r) = 0"


"r = \dfrac{9500}{300} = 31.7 cm"


"x = 30 + 31.7 = 61.7 \text{ cm mark }"